OpenStudy (anonymous):

What is the area of a circle with a radius of 7?

OpenStudy (anonymous):

oh it is 49

OpenStudy (anonymous):

area of a cirlce = pi*r^2 area = 3.14*7*7

OpenStudy (anonymous):

It's wrong @Converse_Luv , i think u forgot to add the value of pi.

OpenStudy (anonymous):

area = 153.86 cm or m ^2

OpenStudy (anonymous):

Thanks a bunch i just did the first step

OpenStudy (anonymous):

its ok, it happens sometimes with me also!

OpenStudy (anonymous):

thanks